OpenStudy (anonymous):

evaluate the following limit:

OpenStudy (anonymous):

\[\lim_{x\rightarrow 9}\frac{3-\sqrt{x}}{x-9} \]

OpenStudy (anonymous):

the answer should be -1/6

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

nevermind, I got this one. Sorry

OpenStudy (campbell_st):

factorise the denimonator and re write the numerator \[\lim_{x \rightarrow 9} \frac{-(\sqrt{x} - 3)}{(\sqrt{x} -3)(\sqrt{x} + 3)}\] eliminate the common factor gives so evaluate \[\lim_{x \rightarrow 9} \frac{-1}{\sqrt{x} + 3}\]